OpenStudy (anonymous):

Eliminate the parameter. x = 7t, y = t + 4 i dont get this at all

OpenStudy (anonymous):

solve x=7t for t. This gives you t=x/7. Then just plug this value of t into the equation for y and get your answer.

jimthompson5910 (jim_thompson5910):

or you can solve y = t + 4 for t to get t = y - 4 now plug this into x = 7t, then solve for y

OpenStudy (anonymous):

so it would be y = (x/7)+4?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

so then would that be the answer?

OpenStudy (anonymous):

yup. definitely. you're done.

OpenStudy (anonymous):

sweet, thanks!

OpenStudy (anonymous):

you're welcome